# Lattice tuning of charge and spin transport in $\beta_{12}$-borophene nanoribbons

### v1

@@ Abstract
Lattice vibrations critically shape charge and spin transport by governing carrier scattering, spin–charge interactions, and spectral redistribution in nanostructures. In this study, we investigate how electron–phonon coupling (EPC) and structural configurations intertwine in magnetic and nonmagnetic $\beta_{12}$-borophene nanoribbons (BNRs). Using a tight-binding framework with site-dependent hopping parameters extracted from ab initio calculations and incorporating phonons within the Holstein model, we compute phonon-renormalized Green’s functions and transport currents via the Landauer–Büttiker formalism. We find that spin-dependent EPC enhances spin-dependent current in magnetic zigzag (ZZ) ribbons, driven by phonon-induced inelastic scattering and spin-selective band renormalization. Additionally, we observe an enhancement of charge transport current in the nonmagnetic configurations of $\beta_{12}$-BNRs. Structural variations further induce anisotropic EPC effects, significantly reshaping charge and spin transport. These insights establish EPC as a powerful design lever for optimizing borophene-based logic devices through tailored edge engineering.
@@ I Introduction
Understanding the role of phonons in charge and spin transport within ultrathin, high-mobility borophene ribbons is essential for the engineering of next-generation electronic and thermal management applications. Building upon these developments, we pose the question: How electron-phonon interactions incorporated by site-dependent electronic hopping energies can influence the quantum transport characteristics of $\beta_{12}$-BNRs?
To systematically explore phonon-mediated transport phenomena, we employ a tight-binding Hamiltonian approach and the nonequilibrium Green’s function method combined with the Landauer-Büttiker formalism [35] and the Holstein model [36, 37, 38]. We investigate both magnetic and non-magnetic BNRs classified by their edge configurations. Moreover, we calculate site-dependent hopping parameters through density functional theory (DFT) simulations and Wannier orbital constructions, enabling a realistic model.
@@ II Methodology
Additionally, maximally localized Wannier functions were constructed using the Wannier90 code [40, 41], interfaced with the Vienna Ab initio Simulation Package [42, 43, 44]. The resulting Wannier Hamiltonians were used as the foundation for detailed investigations of electronic transport and EPC-induced scattering processes.
@@ II.3 Holstein phonon self energy
The fully interacting electronic Green’s function incorporates phonon effects through the phonon self-energy within the NEGF framework. The resulting Green’s function $G_{\gamma\eta}$ is then related to the spectral function $A_{\gamma\eta}(\mathbf{k},\omega)$ in the usual way.  [51]:
@@ III Results and Discussion
In our calculations, we distinguish between two scenarios: a purely coherent regime without phonon effects and a regime with phonons explicitly introduced through Holstein phonons. In the absence of EPC, the lattice is idealized as static, and all carriers propagate coherently without inelastic scattering. Conversely, with EPC, we include site-localized phonons coupled to the electronic degrees of freedom via matrix elements and hopping parameters extracted from our DFT calculations. The numerical results presented in this work are specifically performed at $T=300$ K, where the finite thermal phonon population significantly enhances electron–phonon scattering effects compared to the zero-temperature limit.
@@ III.1 EPC effects on electronic properties of $\beta_{12}$-BNRs
The observation of spin polarization solely in the electronic states of zigzag (ZZ)-terminated $\beta_{12}$-BNRs, as depicted in Fig.8, justifies study of their spin-dependent transport phenomena. To explore these edge-induced magnetism in $\beta_{12}$-BNRs, we have computed site-dependent hopping parameters for each spin channel, detailed in Tables S2, S4, and S6 of the SM [48], corresponding to ribbons with edge widths of 4, 5, and 6 unit cells, respectively. Additionally, site-dependent on-site potentials were computed and are summarized in Tables S3, S5, and S7 of the SM [48] for the same ribbon widths. In addition to spin-dependent hopping parameters, we explicitly compute spin-dependent EPC matrix elements within polarized scenarios, capturing essential spin-dependent features governing the transport behavior of $\beta_{12}$-BNRs. Figure 3 illustrates the interplay among edge geometry (zigzag and line-line), spin polarization, and electron–EPC in a 4-unit-cell-wide $\beta_{12}$-BNR. The LDOS plots correspond to the first supercell, where atom indices match those shown in Fig. 8 in the Appendix.
@@
Although past research has covered borophene’s electronic, magnetic, and phononic properties in various polymorphs [53, 54, 55, 45, 31, 30, 56, 57], most of it either focused on pure 2D sheets or used simplified ribbon setups—often overlooking the detailed impact of EPCs on transport. In this work, we present a more detailed analysis of $\beta_{12}$-BNRs. We combine a localized-orbital tight-binding approach, Holstein phonons, and a two-probe NEGF method to account for edge-induced magnetism, inelastic phonon scattering, and a wide range of ribbon widths and lengths. Our findings reveal spin-dependent renormalization of edge-localized states in ZZ $\beta_{12}$-BNRs, while also showing how phonon scattering affects transport in line–line, AC-AC, and AC–line edge geometries. Furthermore, our device-scale simulations offer practical guidance for engineering borophene-based nanoelectronic and spintronic devices by tailoring edge structure and thermal conditions.
Figure LABEL:fig:LDOS_4ZZ illustrates in detail how edge geometry (zigzag and line–line), spin polarization, and EPC collectively shape the electronic states of a 4-unit-cell-wide $\beta_{12}$. Figures (a–c) correspond to the spin-up ($\uparrow$) channel of the ZZ edge configuration. Without EPC [panel (a)], pronounced localized electronic states emerge near $0.35\,\mathrm{eV}$, strongly concentrated at atoms 17–20 located at the upper edge. This intense localization originates from quantum confinement and under-coordination inherent to zigzag edges, consistent with previous theoretical findings [58, 59]. Upon inclusion of EPC [panel (b)], these edge states broaden substantially and shift upward indicating phonon-mediated energy renormalization. The total DOS in panel (c), computed across the entire four-supercell scattering region, reflects these modifications, exhibiting broadened peaks around $-0.50\,\mathrm{eV}$ and $+0.35\,\mathrm{eV}$. Figures (d–f) describe the spin-down ($\downarrow$) channel of the same zigzag geometry. Here, localized states appear around $-0.40\,\mathrm{eV}$ without EPC [panel (d)], predominantly localized at atoms 6–8. With EPC activated [panel (e)], these localized states are dramatically broadened and experience an approximately $30\,\mathrm{meV}$ downward shift, nearly doubling their spectral width. The total DOS [panel (f)] captures these spectral changes, exhibiting distinctly broader and shifted peaks. Furthermore, increasing the borophene nanoribbon width from 5 to 6 unit cells alters the edge-localized states (as shown in Figure S2 of the SM [48]) increases the number and complexity of edge-localized states, introducing additional phonon modes and scattering pathways that can reinforce or suppress specific features in the DOS. This interrelation between edge geometry, phononic excitations, and spin polarization makes low-dimensional borophene systems sensitive to EPC. Overall, the distinct crystallographic characteristics of ZZ edges in $\beta_{12}$-BNRs centered on the undercoordination of edge atoms—amplify localized states and augment their susceptibility to phonon interactions, rendering EPC a crucial factor in accurately describing the electronic and spin-dependent properties of these systems.
To examine the transport properties of a non-magnetic $\beta_{12}$-BNR, Figure 3(g-i) depicts the LDOS and total DOS for a four-unit-cell-wide ribbon with line-line edges. Calculations are shown (g) without EPC, (h) with EPC, and (i) the integrated DOS for the entire ribbon. Only the first supercell is highlighted in the LDOS panels to emphasize where edge-localized states appear (atom indices correspond to Fig. 8 in the Appendix.). In the line-line configuration, the uniform edge structure produces sharp peaks at discrete energies, indicative of strong quantum confinement. Edge-localized states, intrinsic to the finite ribbon width, are highly sensitive to electron-phonon interactions, which significantly affect the transport characteristics. For additional configurations (AC-AC and AC-line configurations), see Figs. S3, S4 and S5 of SM [48].
III.2 Spin-polarized charge transport in $\beta_{12}$-BNRs with EPC
@@
Figure 4(a) illustrates the two-probe device structure employed in our nonequilibrium Green’s function calculations of electron transport through a zigzag-edge $\beta_{12}$-BNR. The scattering region consists of four supercells, each four fundamental unit cells wide; EPC is explicitly included only within this finite central region, while the left and right electrodes remain ballistic, acting as ideal reservoirs. Panels (b) and (c) show spin-resolved current–voltage ($I$$V_{b}$) characteristics and their associated differential conductances (${\rm d}I/{\rm d}V_{b}$), respectively, comparing scenarios with (solid lines) and without EPC (dashed lines). Inclusion of EPC increases the currents in both spin channels throughout the applied bias range. Figure 4(d) decomposes this enhanced EPC current into coherent ($I_{0}$) and phonon-assisted ($I_{\mathrm{ph}}$) contributions, indicating that the net current increase is dominated by phonon-mediated inelastic processes. As illustrated in Fig. 3, the higher spin-down current originates from the larger and more continuous DOS around the Fermi level in the spin-down channel, offering more accessible electronic states for conduction. Consequently, the coherent contribution slightly decreases due to enhanced electron scattering. The bias-dependent differential conductance curves shown in Figs. 3(c) and (e) remain nearly constant under ballistic conditions, consistent with the approximately linear dispersion characteristic of metallic ribbons near the Fermi level. Inclusion of EPC increases the overall differential conductance and introduces a subtle upward slope at higher biases, especially in the spin-up channel. This smooth enhancement is attributed to additional inelastic transmission channels that become accessible once electrons can exchange energy through interactions with phonons.
@@
[…]
Figures 5(e-h) illustrate how transport characteristics vary with ribbon width at fixed length. In the ballistic regime shown in Fig. 5(e), the current increases nearly linearly with applied bias, displaying a non-monotonic dependence on ribbon width at $_{b}=1,\text{V}$ V, total currents for widths $N_{W}=4,5,6$ are approximately 484848, 353535, and $61,\mu\text{A}$, respectively. Notably, the intermediate-width ribbon ($N_{W}=5$) exhibits a pronounced reduction in current relative to the even-width ribbons. This phenomenon is attributed to significant hybridization of edge-localized states, particularly affecting the spin-down ($\downarrow$) channel and consequently reducing the number of available low-energy conduction pathways. The ballistic differential conductance [Fig. 5(f)] remains largely bias-independent, reflecting stable conduction channel populations. Even-width ribbons ($N_{W}=4,6$) display notably higher conductance in the spin-down channel compared to spin-up, due to clearly defined, non-hybridized edge states positioned close to the Fermi level. In contrast, the odd-width ribbon ($N_{W}=5$) exhibits markedly lower conductances in both spin orientations, consistent with partial hybridization-induced suppression of the edge states. Introducing EPC significantly enhances both the total current and differential conductance [Figs. 5(g),(h)], contrary to the typical expectation of transport suppression by phonon scattering. This enhancement arises due to phonon-assisted inelastic tunneling processes, where electrons emit or absorb phonons, effectively overcoming hybridization gaps or small conduction barriers and thereby increasing the accessible conduction pathways. Remarkably, the intermediate-width ribbon ($N_{W}=5$) experiences the strongest EPC-driven improvement, largely mitigating the hybridization-induced suppression observed in the ballistic regime and restoring a nearly monotonic width-dependent trend. The EPC-driven enhancement is further evidenced in the differential conductance profiles [Fig. 5(h)], which shift upward significantly compared to their ballistic counterparts. Rather than merely broadening the quantized conductance steps, phonons introduce additional conduction channels, particularly visible at higher biases. The spin-up channel, initially weaker in the ballistic limit, is notably strengthened under EPC, altering the relative spin conductance hierarchy. This phonon-mediated adjustment indicates an effective means to modulate spin-dependent transport in $\beta_{12}$-BNRs, emphasizing the critical interplay between structural (width-dependent edge-state hybridization) and vibrational (phonon-assisted tunneling) properties in shaping the transport characteristics of these nanostructures.

### v2

@@ Abstract
$\beta_{12}$-borophene nanoribbons (BNRs) exhibit magnetic zigzag edges, while other edge configurations are nonmagnetic. However, when the source, central, and drain regions of a logic device are all composed of zigzag BNRs (ZBNRs), the resulting spin polarization remains weak, unless a high voltage is applied. In this work, we demonstrate that lattice vibrations—introduced for example, via a thermal bath coupled to the central BNR—can enhance spin polarization in ZBNRs. This enhancement manifests as marked changes in the current-voltage characteristics, enabling direct experimental probing. In contrast, nonmagnetic edge configurations exhibit phonon-enhanced charge transport. We employ a tight-binding approach augmented with local electron-phonon interactions described by the Holstein model, and compute the phonon-renormalized Green’s functions and transport currents using the Landauer-Büttiker formalism. The mechanism is supported by analyzing both spinless and spinful electronic dispersions and the corresponding density of states. Compared to the phonon-free edges, structural distortions lead to anisotropic electron-phonon couplings, which significantly modify both charge and spin transport. These results position phonon as an effective tuning parameter for optimizing borophene-based logic devices via engineered edge configurations.
@@ I Introduction
Understanding the role of phonons in charge and spin transport within ultrathin, high-mobility borophene ribbons is also essential for the engineering of next-generation electronic and thermal management applications. Building upon these developments, we pose the question: How do phonons affect the quantum transport properties of $\beta_{12}$-borophene?
To systematically explore phonon-mediated transport phenomena, we employ a tight-binding Hamiltonian approach and the nonequilibrium Green’s function method combined with the Landauer-Büttiker formalism [35] and the Holstein model [36, 37, 38]. We investigate both magnetic and non-magnetic BNRs classified by their edge configurations.
@@ II.3 Holstein phonon self energy
The fully interacting electronic Green’s function incorporates phonon effects through the phonon self-energy within the NEGF framework. The resulting Green’s function $G_{\gamma\eta}$ is then related to the spectral function $A_{\gamma\eta}(\mathbf{k},\omega)$ in the usual way [46]:
@@ III Results and Discussion
In our calculations, we distinguish between two scenarios: a purely coherent regime without phonon effects and a regime with phonons explicitly introduced through Holstein phonons. In the absence of EPC, the lattice is idealized as static, and all carriers propagate coherently without inelastic scattering. Conversely, with EPC, we include site-localized phonons coupled to the electronic degrees of freedom via matrix elements extracted from our DFT calculations. The numerical results presented in this work are specifically performed at $T=300$ K, where the finite thermal phonon population significantly enhances electron–phonon scattering effects compared to the zero-temperature limit.
@@ III.1 EPC effects on electronic properties of $\beta_{12}$-BNRs
Although past research has covered borophene’s electronic, magnetic, and phononic properties in various polymorphs [48, 49, 50, 40, 31, 30, 51, 52], most of it either focused on pure 2D sheets or used simplified ribbon setups—often missing the detailed impact of EPCs on transport. In this work, we present a more detailed analysis of $\beta_{12}$-BNRs. We combine a localized-orbital tight-binding approach, Holstein phonons, and a two-probe NEGF method to account for edge-induced magnetism, inelastic phonon scattering, and a wide range of ribbon widths and lengths. We explicitly compute spin-dependent EPC matrix elements within polarized scenarios, capturing essential spin-dependent features governing the transport behavior of $\beta_{12}$-BNRs. The observation of spin polarization solely in the electronic states of zigzag (ZZ)-terminated $\beta_{12}$-BNRs, as depicted in Fig.8, justifies study of their spin-dependent transport phenomena. Our findings reveal spin-dependent renormalization of edge-localized states in ZZ $\beta_{12}$-BNRs, while also showing how phonon scattering affects transport in line–line, AC-AC, and AC–line edge geometries. Furthermore, our device-scale simulations offer practical guidance for engineering borophene-based nanoelectronic and spintronic devices by tailoring edge structure and thermal conditions.
Figure 3 illustrates in detail how edge geometry, spin polarization, and EPC collectively shape the electronic states in a 4-unit-cell-wide $\beta_{12}$-BNR. Panels (a–f) represent the zigzag-edge configuration, while panels (g–i) correspond to the line–line edge (non-magnetic) configuration. The left panels (a,d,g) show the layer-resolved local density of states (LDOS) without EPC, middle panels (b,e,h) include EPC, and right panels (c,f,i) compare the total DOS, integrated over the entire scattering region, with (red curves) and without EPC (blue curves). The atomic indices and their corresponding positions for each configuration are explicitly shown in Fig. 8.
@@
In the spin-down channel for the zigzag-edge configuration without EPC [panel (d)], prominent states emerge distinctly at energies around $E\approx-0.18\,\mathrm{eV}$ and $E\approx 0.28\,\mathrm{eV}$, again showing strong intensity localized predominantly but not exclusively at edge atoms. Crucially, these states display substantial LDOS contributions from inner atomic rows as well, reflecting their partially delocalized character. Incorporation of EPC [panel (e)] significantly alters this spectral distribution. The states around $-0.18\,\mathrm{eV}$ and $0.28\,\mathrm{eV}$ become broadened, with greatly reduced intensity, consistent with electron–phonon-induced lifetime shortening and enhanced inelastic scattering. Furthermore, EPC affects LDOS intensity in interior regions, indicative of increased hybridization between edge-derived and bulk-derived states. The total DOS [panel (f)] quantitatively confirms these EPC-induced alterations, displaying pronounced broadening and intensity reduction of peaks around $-0.18\,\mathrm{eV}$ and $0.28\,\mathrm{eV}$, alongside smoother spectral distributions, underscoring substantial phonon-mediated spectral reconfiguration.
Furthermore, increasing the borophene nanoribbon width from 5 to 6 unit cells alters the edge-localized states (as shown in Figs. S2 of the SM [43]) increases the number and complexity of edge-localized states, introducing additional phonon modes and scattering pathways that can reinforce or suppress specific features in the DOS. This interrelation between edge geometry, phononic excitations, and spin polarization makes low-dimensional borophene systems sensitive to EPC. Overall, the distinct crystallographic characteristics of ZZ edges in $\beta_{12}$-BNRs centered on the undercoordination of edge atoms—amplify localized states and augment their susceptibility to phonon interactions, rendering EPC a crucial factor in accurately describing the electronic and spin-dependent properties of these systems.
To examine the transport properties of a non-magnetic $\beta_{12}$-BNR, Fig. 3(g-i) depicts the LDOS and total DOS for a four-unit-cell-wide ribbon with line-line edges. Calculations are shown (g) without EPC, (h) with EPC, and (i) the integrated DOS for the entire ribbon. Only the first supercell is highlighted in the LDOS panels to emphasize where edge-localized states appear (atom indices correspond to Fig. 8 in the Appendix.). In the line-line configuration, the uniform edge structure produces sharp peaks at discrete energies, indicative of strong quantum confinement. Edge-localized states, intrinsic to the finite ribbon width, are highly sensitive to electron-phonon interactions, which significantly affect the transport characteristics. For additional configurations (AC-AC and AC-line configurations), see Figs. S3, S4 and S5 of SM [43].
III.2 Spin transport in magnetic $\beta_{12}$-BNRs with EPC
As we observe spin polarization exclusively in $\beta_{12}$-BNRs with ZZ edges, here we investigate how EPC influences the spin transport in these systems. Our findings demonstrate the key role of EPC in modulating the spin-dependent electronic structure of ZZ $\beta_{12}$-BNRs, shedding light on the interplay between edge morphology, spin order, and electron-phonon interactions. The spin polarization itself originates from introducing a rigid exchange splitting ($\Delta=0.20\,\text{eV}$ [22, 24]), applied uniformly to the spin-degenerate non-magnetic Hamiltonian, as detailed explicitly in Eq. (S1) of the Supplemental Material. In our transport simulations, the magnetic ZZ $\beta_{12}$-BNR serves as the central scattering region, sandwiched between non-magnetic ZZ $\beta_{12}$-BNR electrodes, ensuring that spin polarization emerges solely from the constructed spin-dependent Hamiltonian of the scattering region.
@@ III.2 Spin transport in magnetic $\beta_{12}$-BNRs with EPC
Figure 4(a) illustrates the two-probe device configuration employed in our NEGF calculations of electron transport through a ZZ $\beta_{12}$-BNRs. The central scattering region consists of four supercells, each composed of four fundamental unit cells along its width; EPC is explicitly included only within this finite region, while the left and right electrodes are modeled as semi-infinite ballistic reservoirs. Panels (b) and (c) present the spin-resolved current–voltage ($I$$V_{b}$) characteristics and their corresponding differential conductances (${\rm d}I/{\rm d}V_{b}$), respectively, comparing results obtained with EPC (solid lines) and without EPC (dashed lines). Under coherent conditions, panel (b) shows near-linear current–voltage relationships, reflecting the intrinsic metallicity of these ribbons. At low biases, the currents and differential conductances for both spin channels nearly overlap, indicating negligible spin polarization in this regime. However, as the bias increases, distinct spin polarization emerges and progressively intensifies, a phenomenon clearly discernible from the divergence of the spin-up and spin-down branches in panel (b) and, more evidently, from the differential conductance curves in panel (c). Notably, in the presence of EPC, spin polarization appears at significantly lower bias voltages compared to the ballistic case. Furthermore, the inclusion of EPC consistently enhances the total current in both spin channels across the entire bias range. To elucidate the origin of this current enhancement, Fig. 4(d) decomposes the total EPC-influenced current into coherent ($I_{0}$) and phonon-assisted ($I_{\mathrm{ph}}$) contributions, revealing that the net current increase predominantly arises from phonon-mediated inelastic processes. Interestingly, the coherent component ($I_{0}$) remains essentially unaffected by EPC, confirming that elastic transmission channels are only weakly perturbed by electron–phonon interactions. In contrast, the phonon-assisted current ($I_{\mathrm{ph}}$) grows substantially with bias, highlighting the pivotal role of inelastic electron–phonon scattering in activating additional conduction pathways. A detailed analysis of the differential conductance is further provided in Fig. 3(e), where the total differential conductance is explicitly resolved into its coherent and phonon-assisted components. Under purely ballistic conditions (without EPC), the differential conductance already displays significant bias dependence, rising gradually from about $50\,\mu$A V-1 at low bias and exhibiting a broad maximum around $V_{b}\approx 2$ V, followed by a slight reduction at higher voltages. This non-monotonic behavior directly reflects the intrinsic electronic band structure of the metallic ribbon, influenced by sub-band alignments and van Hove singularities in the density of states near the Fermi energy. Upon introducing EPC, pronounced peaks emerge sharply in the differential conductance around $V_{b}\approx 1.1$ and 1.71.71.7 V, corresponding precisely to resonant inelastic scattering events mediated by specific phonon modes. At these characteristic biases, electrons efficiently exchange discrete energy quanta with phonons, accessing otherwise inaccessible electronic states and dramatically boosting the conductance. Consequently, the net effect of EPC manifests as a substantial overall increase in differential conductance, dominated by phonon-mediated inelastic contributions.
@@
Figure 5 compares the current–voltage and differential conductance characteristics of ZZ $\beta_{12}$-BNRs under different transport regimes. Panels (a)–(d) focus on a ribbon with width $NW=4$ and lengths $NL=1,4,7,10$, whereas panels (e)–(h) examine a fixed length across widths $NW=4,5,6$. Solid (dotted) lines indicate spin-up (spin-down) transport, respectively. In all cases, we distinguish between no inelastic scattering and EPC-influenced transport, as well as spin-up and spin-down channels. Under coherent conditions, panel (a) the current magnitude systematically declines as the ribbon length increases, consistent with exponential decay due to coherent tunneling limitations within the extended scattering region. As highlighted by the inset of panel (a), this length-induced suppression remains significant even at high biases. The associated differential conductance curves in panel (b) exhibit pronounced multi-hump features: initially rising near $V_{b}\sim 0.7$ V, subsequently decreasing around 111 V, followed by secondary peaks near 1.31.31.3 V and finally leveling off beyond 2.22.22.2 V. These oscillations originate from sub-band crossings and van-Hove singularities in the ribbon’s density of states and diminish in amplitude for longer ribbons due to reduced coherent transmission. Moreover, spin-resolved differences become noticeable at biases above 1$\sim$ 1∼ 1 V, with spin-down conductance consistently exceeding spin-up conductance, thereby revealing a clear bias-induced spin polarization amplified with ribbon length. Introducing EPC alters this scenario, as demonstrated in panels (c) and (d). Panel (c) reveals that EPC enhances the current across the entire bias range for all ribbon lengths, notably reversing the length dependence observed in panel (a): the longest ribbon ($NL=10$) now exhibits the highest current. This reversal arises from the dominance of phonon-mediated inelastic transmission, whose contribution scales approximately linearly with increasing ribbon length and thereby surpasses the coherent tunneling limitations. Correspondingly, the differential conductance curves in panel (d) undergo substantial reshaping: EPC elevates the conductance from approximately $25\,\mu\mathrm{A\,V}^{-1}$ (without EPC) to nearly $30\,\mu\mathrm{A\,V}^{-1}$ for $NL10$ at zero bias, broadens and smooths the ballistic humps into distinct plateaus, and introduces additional resonant peaks around biases of 111 V and 1.51.51.5 V. These features reflect bias-dependent activation of specific optical-phonon modes that provide efficient channels for electron energy exchange, thus unlocking otherwise inaccessible conduction pathways.
Figures 5(e)–(h) systematically examine how varying ribbon width ($NW=4$, 555, 666) at a fixed length affects the transport characteristics in zigzag-edge $\beta_{12}$-BNRs. Panels (e) and (f) represent coherent transport (w/o EPC), whereas panels (g) and (h) include EPC (w/ EPC). In the ballistic limit [panel (e)], the $I$$V_{b}$ characteristics are nearly linear, and the magnitude of the current systematically increases with ribbon width, reflecting a greater number of transverse conduction modes available in wider ribbons. Spin-down currents consistently surpass spin-up currents at higher biases ($V_{b}\gtrsim 2$ V), indicating a clear bias-induced spin polarization arising from width-dependent spin-split states localized at the ribbon edges. Corresponding ballistic differential conductances in panel (f) exhibit distinct multi-peak structures, associated with sub-band crossings and van-Hove singularities in the electronic density of states. These peaks become increasingly pronounced with ribbon width, owing to enhanced transmission through additional conduction pathways. Moreover, spin-down conductance consistently exceeds spin-up across most bias voltages, highlighting intrinsic spin-dependent asymmetry driven by ribbon geometry and edge states. Upon inclusion of EPC, transport properties dramatically evolve [panels (g) and (h)]. Panel (g) reveals a pronounced EPC-induced enhancement in total current for all ribbon widths. Interestingly, this enhancement is non-monotonic with respect to width, with the intermediate-width ribbon ($NW=5$, green lines) exhibiting the highest total current, surpassing both narrower ($NW=4$) and wider ($NW=6$) ribbons. Such behavior results from competing effects: although wider ribbons offer more scattering sites and conduction channels for phonon-assisted transport, they simultaneously experience enhanced backscattering from increased phonon-induced disorder, thus partially limiting their conductance advantage. Crucially, the spin-dependent current behavior under EPC is notably modified compared to the ballistic regime. In particular, the spin-up current surpasses the spin-down current for the wider ribbons ($NW=5$ and $NW=6$). This remarkable reversal in spin conductance ordering arises due to selective phonon-mediated enhancement of transmission channels predominantly associated with spin-up edge states. Electron–phonon scattering preferentially couples spin-up electrons to previously inaccessible conduction pathways, effectively redistributing spectral weight toward spin-up channels and overcoming their ballistic suppression relative to spin-down channels. This phenomenon is further corroborated by differential conductance curves in panel (h), where EPC-induced resonance peaks with maximum values reaching around $280\,\mu\text{A\,V}^{-1}$ for the $NW=5$ ribbon. These peaks directly result from resonant electron–phonon processes involving specific optical phonon modes, facilitating spin-dependent energy exchange and opening new phonon-mediated conduction channels. Consequently, EPC not only boosts overall conduction efficiency but also provides an effective mechanism for altering spin polarization, particularly by strengthening spin-up conductance beyond spin-down levels in wider ribbons. Thus, these results illustrate how the delicate interplay between electronic structure, phononic interactions, and ribbon geometry influences the transport characteristics, enabling precise control over spin-dependent currents in zigzag-edge BNRs.